# Characteristics of Dark Current from a S-band RF Gun with Exchangeable Photocathode System

### v3

@@ IV RESULTS AND DISCUSSION
For the simulated images shown in Fig. 5, the colour map (middle row) used is the charge per 2D bin, to compare with the intensity colour map of the screen images. The comparison of similar features in screen measurements and simulations shown in Fig. 5 indicate that dark current originating from field emission from the rims in the cathode region give rise to the distinctive ‘ring’ features observed. This is clarified by the colour coded comparison of the separated contributions from only the rims (Fig. 5 bottom row). In the first solenoid configuration this leads to a bright ring of roughly 25 mm diameter originating from the backplate rim and a dimmer ring of roughly 5 mm originating from the cathode rim. In the second solenoid configuration the bright central spot originates from the cathode rim with a larger ‘halo’ originating from the backplate rim. In the final solenoid configuration a bright ring of roughly 7 mm originates from the cathode rim. There may be several factors involved in the relative brightness of rings due to the entire transport of the dark current from source to screen, but more intense emission from the rim surfaces compared to the flat central cathode surface may be expected due to electric field enhancement. The surface roughness on the rims may be expected to be greater than the flat cathode surface and microscopic field enhancement on the (greater) geometric surface imperfections of the rims would generate more field emission. A more complete formulation of the FN model [17] defines surface imperfections by a field enhancement factor $\beta$ which scales the effective electric field experienced on the surface. There is some uncertainty in using this model to extract quantitative values of $\beta$ for RF guns [17], but the model predicts a strong dependence of the amount of emission on surface quality, which supports the reasoning above. In addition, the non-enhanced electric field strength varies over the cathode region and reaches maxima on the rims (see Fig. 3.) which again implies an increased contribution to dark current of the rims.

### v4

@@ IV RESULTS AND DISCUSSION
For the simulated images shown in Fig. 5, the colour map (middle row) used is the charge per 2D bin, to compare with the intensity colour map of the screen images. The comparison of similar features in screen measurements and simulations shown in Fig. 5 indicate that dark current originating from field emission from the rims in the cathode region give rise to the distinctive ‘ring’ features. This is clarified by the colour coded comparison of the separated contributions from only the rims (Fig. 5 bottom row). In the first solenoid configuration this leads to a large bright ring originating from the backplate rim, and a dimmer smaller ring originating from the cathode rim. In the second solenoid configuration the bright central spot originates from the cathode rim with a larger ‘halo’ originating from the backplate rim. In the final solenoid configuration a bright ring of originates from the cathode rim. There may be several factors involved in the relative brightness of rings due to the entire transport of the dark current from source to screen, but more intense emission from the rim surfaces compared to the flat central cathode surface is expected due to electric field enhancement. The surface roughness on the rims is expected to be greater than the flat cathode surface, thus field enhancement on the rims would generate more field emission (a more complete formulation of the FN model [17] includes parameterisation of roughness in a field enhancement factor $\beta$, and predicts a strong dependence of the amount of emission on surface quality). In addition, the non-enhanced electric field strength varies over the cathode region and reaches maxima on the cathode rim (see Fig. 3.) which again implies an increased contribution to dark current from the cathode rim.